# Twin Deformable Point Convolutions for Point Cloud Semantic Segmentation in Remote Sensing Scenes

### v1

@@ Abstract
Thanks to the application of deep learning technology in point cloud processing of the remote sensing field, point cloud segmentation has become a research hotspot in recent years, which can be applied to real-world 3D, smart cities, and other fields. Although existing solutions have made unprecedented progress, they ignore the inherent characteristics of point clouds in remote sensing fields that are strictly arranged according to latitude, longitude, and altitude, which brings great convenience to the segmentation of point clouds in remote sensing fields. To consider this property cleverly, we propose novel convolution operators, termed Twin Deformable point Convolutions (TDConvs), which aim to achieve adaptive feature learning by learning deformable sampling points in the latitude-longitude plane and altitude direction, respectively. First, to model the characteristics of the latitude-longitude plane, we propose a Cylinder-wise Deformable point Convolution (CyDConv) operator, which generates a two-dimensional cylinder map by constructing a cylinder-like grid in the latitude-longitude direction, and then performs adaptive feature sampling on the cylinder map by deformable offset learning. Furthermore, to better integrate the features of the latitude-longitude plane and the spatial geometric features, we perform a multi-scale fusion of the extracted latitude-longitude features and spatial geometric features, and realize it through the aggregation of adjacent point features of different scales. In addition, a Sphere-wise Deformable point Convolution (SpDConv) operator is introduced to adaptively offset the sampling points in three-dimensional space by constructing a sphere grid structure, aiming at modeling the characteristics in the altitude direction. Experiments on existing popular benchmarks conclude that our TDConvs achieve the best segmentation performance, surpassing the existing state-of-the-art methods.
Index Terms:
Remote Sensing, Airborne Laser Scanning, Point Cloud, Semantic Segmentation, Geographic Information Modeling, Deformable Convolution, Deep Learning
I Introduction
@@
In recent years, with the rapid advancement of deep learning technology and the explosive growth of remote sensing satellite data, the interpretation methods of remote sensing data have achieved unprecedented development in the fields of classification[34], segmentation[39], object detection[43], height estimation[48], and 3D reconstruction[53]. Especially the research on 3D remote sensing data has become increasingly popular. As the main sensor for collecting point cloud data in the field of remote sensing, airborne laser scanning (ALS) technology, has gradually entered the field of vision of researchers. Among them, ALS point cloud segmentation has received a lot of attention.
@@
Figure 1: ALS point clouds are acquired by the aircraft, therefore, these point clouds are strictly arranged according to the latitude and longitude direction and the altitude direction. (a) Latitude and Longitude Distribution. The instance objects in the large-scale point cloud are strictly arranged in fixed latitude and longitude positions. (b) Altitude Distribution. Different instance objects have different altitudes, which is one of the inherent characteristics of large-scale city-level point clouds.
@@
II Related Work
II-A Classical Machine Learning-Based Methods
@@
II-B Point Cloud Segmentation Based on Deep Learning
The rapid development of deep learning has promoted the replacement of manual feature design methods by automatic point cloud feature extraction[10, 3, 6]. Among them, point-based methods are the most favored by researchers among many methods. The point-based methods refer to directly processing the point cloud without any data preprocessing operation, and then achieving the classification and segmentation tasks of the input point set by learning the deep semantic features of the point set. It mainly includes MLP-based methods[6, 5, 11, 33], graph-based methods[9, 54, 12], and point convolution-based methods[8, 57, 7, 3, 17]. MLP-based methods[6, 5, 11, 33] are the earlier deep convolutional neural network methods for processing point cloud data. PointNet[6] is the pioneering work of this type of method. It first uses multiple multi-layer perceptrons with shared weights to learn point-by-point high-dimensional features, and then designs a symmetric function based on the maximum pooling function to solve the disorder problem of point cloud. PointNet++[5] models local relationships through the three processes of sampling, grouping, and feature learning inside the point cloud. Point sets are the discrete structure, and there is a certain connection relationship between the points. This is a natural graph structure data form. Therefore, some researchers have introduced graph structure modeling into point cloud learning and achieved excellent results. Landrieu et al.[18] earlier proposed a point cloud convolutional neural network SPGraph based on graph convolution, which used graph convolution to learn high-level semantic features of the graph structure, and ultimately achieved point-by-point semantic segmentation. Although traditional convolution has been successful in feature extraction of 2D raster image data, it cannot be directly transferred to point cloud processing problems due to the unique properties of point cloud data. Therefore, Li et al.[8] tried to weight the features of the input point set and use a multi-layer perceptron to learn the X transformation to achieve point-by-point weighted summation in the neighborhood. In addition, Li et al[57] proposed a geometric attention network composed of geometry-aware convolution, dense hierarchical architecture, and elevation attention module, and trained in an end-to-end manner to solve the three problems (geometric instances, extreme scale changes, and large differences in elevation) of remote sensing point clouds.
@@ II-B Point Cloud Segmentation Based on Deep Learning
Figure 2: Overall Architecture. Our TDConvs adopt the popular Encoder-Decoder structure as the basic component. Each feature extraction layer in the encoder part is the proposed Cylinder-wise Deformable point Convolution (CyDConv), and each layer in the decoder part is a feature forward propagation layer. The encoder and decoder are connected as feature aggregation layers with our proposed Sphere-wise Deformable point Convolution (SpDConv) at the highest resolution layer. The dimension of input point sets is $\rm N\times M$, where $\rm M$ is the dimension of point features. The dimension of output point sets is $\rm N\times C$, where $\rm C$ is the number of output categories.
II-C Geospatial Modeling in Point Cloud Segmentation
@@
III Proposed Method
III-A Method Overview
@@
As shown in Fig. 2, the flowchart of our entire framework is given. Given the sampled point set, the encoder and decoder are executed sequentially. In addition, there is a skip-connection layer between each layer of the encoder and decoder for feature fusion. We embed the learning of longitude and latitude information into Cylinder-wise Deformable point Convolution (CyDConv), and the learning of altitude information into Sphere-wise Deformable point Convolution (SpDConv). CyDConv is performed in each feature extraction layer of the encoder, and SpDConv is performed in the highest resolution feature aggregation layer. The decoder is rooted in the feature propagation, which is proposed in PointNet++[5].
III-B Cylinder-wise Deformable Point Convolution
@@
Cylindricization and Map Generation. Specifically, given the input point set $\mathbf{P}\in\mathbb{R}^{\rm N\times 3}$ and its feature $\mathbf{F}\in\mathbb{R}^{\rm N\times M}$, we first normalize the coordinates of the input point set $\mathbf{P}$, and then convert the point cloud into a cylinder structure at the top view, called cylindricization. Mathematically, we first construct the center points through the grid arrangement as shown in the bottom of Fig. 3, which matches each cylinder of the top view, as
@@
Figure 3: The architecture diagram of our proposed Cylinder-wise Deformable point Convolution (CyDConv). ‘LBR’ means the Linear-BN-ReLU operation. The parts from top to bottom in the figure represent Multi-Nearest Neighbor Feature Learning (MFL), Map Feature Sampling and Aggregation, Map Offset Learning, and Cylindricization and Map Generation processes, respectively.
Map Offset Learning. As shown in Fig. 3, to model the relationship within point sets in the latitude and longitude directions, we learn the offset through the point coordinates and point features, as
@@
Multi-Nearest Neighbor Feature Learning (MFL). From the perspective of the cylinder map, we have realized the modeling and learning of latitude and longitude features. For the modeling of geometric features inside point sets, we propose a Multi-nearest neighbor Feature Learning method called MFL. Specifically, we learn point cloud features by constructing neighboring points of different scales and then perform feature aggregation to realize geometric modeling of point cloud features of different scales, as shown in the top of Fig. 3.
Figure 4: The architecture diagram of the proposed Sphere-wise Deformable point Convolution (SpDConv). ‘T’ is a switch for whether to use SpDConv, which determines which connection path between the encoder and decoder uses SpDConv. Only the highest resolution connection path is selected to embed SpDConv in our network.
@@
III-C Sphere-wise Deformable Point Convolution
@@
Spheroidization and Volume Generation. As shown in Fig. 2, our proposed SpDConv is embedded in the feature propagation layer between the encoder and the decoder. Since the highest-resolution point set has the most comprehensive altitude features, our SpDConv only selects the highest-resolution feature aggregation layer for embedding.
From Fig. 4, given the highest resolution point set $\mathbf{P}\in\mathbb{R}^{\rm N\times 3}$ and point feature $\mathbf{F}\in\mathbb{R}^{\rm N\times M}$. We first choose whether to close switch ‘T’. When switch ‘T’ is closed, SpDConv starts to perform operations. If the switch ‘T’ is closed, unlike CyDConv, we first build a three-dimensional grid structure. The size of the grid structure is $\rm H_{s}$,$\rm W_{s}$, and $\rm Z_{s}$, which is expressed by the formula
@@ III-D Loss Function
IV Experiments
IV-A Experimental Setting
IV-A1 Datasets
We selected the ISPRS Vaihingen 3D[20, 2, 26] and LASDU[40] datasets for experimental verification.
@@ IV-A1 Datasets
IV-A2 Implementation Details
In the experiment, we adopted the PyTorch framework to complete our method. Specifically, for the data preprocessing process, we divide the point cloud of the datasets into patches according to a certain distance step (ISPRS Vaihingen 3D: $30m\times 30m$, LASDU: $50m\times 50m$). For the training process, we randomly sample 4096 points from each patch as the input point sets of our TDConvs. In addition, the batch size and learning rate are set as 4 and 0.0002, respectively. We trained for 500 epochs on a single A40 GPU. What’s more, we adopt the Adam optimizer to minimize our target loss function. During testing, we take all points as input and evaluate the output results. The weights of our loss function are set to $\{1.0,2.0,2.0,2.0,2.0\}$ for the ISPRS Vaihingen 3D dataset and $\{1.0,5.0,5.0,5.0,5.0\}$ for LASDU dataset.
TABLE I: Performance of our TDConvs and other methods on the ISPRS Vaihingen 3D dataset. ‘vision’ means vision-based methods. ‘Remote Sensing’ represents remote sensing-based methods. The first 9 columns of the table represent the F1 score of each category. OA: overall accuracy. mF1: mean F1 score. Bold indicates the first performance, and underline indicates the second performance.
@@ IV-A2 Implementation Details
IV-A3 Evaluation Metrics
$\rm Precision=\frac{TP}{TP+FP}$ (20)
$\rm Recall=\frac{TP}{TP+FN}$ (21)
$\rm OA=\frac{TP+TN}{TP+TN+FP+FN}$ (22)
[…]
What’s more, mF1 are the averages of F1 for all categories.
IV-B Performance Analysis
IV-B1 Quantitative Analysis
We present the segmentation performance of our TDConvs on two datasets, ISPRS Vaihingen 3D and LASDU, respectively.
TABLE II: Performance comparison of our TDConvs with other methods on the LASDU dataset.
Bold indicates the first performance and underline indicates the second performance.
@@ IV-B1 Quantitative Analysis
Figure 5: Segmentation results of the proposed TDConvs on the ISPRS Vaihingen 3D test set. The results in the box are local magnifications of the large area.
@@
ISPRS Vaihingen 3D: As shown in Table I, we present the performance comparison of our TDConvs with other point-based methods on the ISPRS Vaihingen 3D dataset. In general, our TDConvs achieve 84.5% on OA and 73.4% on mF1, respectively, surpassing all other remote sensing-based and vision-based methods, which strongly proves the superiority of our TDConvs. Specifically, our TDConvs outperforms the popular RFFS-Net[1] by 2.4% on OA and 1.8% on mF1. OA and mF1 evaluate the overall accuracy and category accuracy of the model, respectively. Obviously, our TDConvs achieves the best performance for both OA and mF1, and significantly outperforms the SOTA method. For specific categories, our TDConvs achieves balanced performance on each category with small variance in performance, and there are no categories with extremely low performance. This is because ALS point clouds are strictly arranged according to latitude, longitude, and altitude. Our TDConvs can achieve balanced segmentation performance. In particular, for categories of car, fence, and facade, our TDConvs achieves the best or second segmentation performance, which strongly demonstrates the superiority of our approach.
Figure 6: Segmentation results of the proposed TDConvs on the LASDU test set. The upper part is the ground truth of the dataset, and the lower part is the segmentation results of the proposed TDConvs.
Figure 7: Comparison of the segmentation results of the proposed TDConvs with other methods on the LASDU test set. The first column is the segmentation results of PointConv[7], the second column is the segmentation results of RFFS-Net[1], and the third column is the segmentation results of the proposed TDConvs.
@@
LASDU: The comparison on the LASDU dataset of our TDConvs and other methods is given in Table II. From the table, it can be concluded that our proposed TDConvs achieves the best segmentation performance, reaching an mF1 score of 77.85%. In addition, on the OA indicator, our method is comparable to the optimal MCFN, but on the mF1 indicator, the TDConvs we proposed is 0.85% higher than MCFN[49], which is a big improvement in ALS point cloud segmentation. Compared with the popular RFFS-Net[1], our proposed TDConvs is 0.49% and 0.16% higher in OA and mF1 indicators respectively, which also reflects the superiority of the method in this paper. Compared with OA, mF1 can better express the performance of point cloud segmentation because it takes into account the performance of each category, which is particularly obvious in datasets with imbalanced categories. Therefore, in the field of segmentation, we focus more on the performance of mF1, and on the LASDU dataset, our TDConvs achieve the best mF1 performance, which strongly proves the superior performance of our TDConvs in the field of point cloud segmentation. As shown in Table II, in specific categories, although our method does not achieve the best segmentation performance, the performance of each category is balanced. This has obvious advantages over IPCONV[50], which has large segmentation differences in categories. This is why our method can achieve the best mean F1 performance.
IV-B2 Qualitative Analysis

### v2

@@ Abstract
[abstract] Thanks to the application of deep learning technology in point cloud processing of the remote sensing field, point cloud segmentation has become a research hotspot in recent years, which can be applied to real-world 3D, smart cities, and other fields. Although existing solutions have made unprecedented progress, they ignore the inherent characteristics of point clouds in remote sensing fields that are strictly arranged according to latitude, longitude, and altitude, which brings great convenience to the segmentation of point clouds in remote sensing fields. To consider this property cleverly, we propose novel convolution operators, termed Twin Deformable point Convolutions (TDConvs), which aim to achieve adaptive feature learning by learning deformable sampling points in the latitude-longitude plane and altitude direction, respectively. First, to model the characteristics of the latitude-longitude plane, we propose a Cylinder-wise Deformable point Convolution (CyDConv) operator, which generates a two-dimensional cylinder map by constructing a cylinder-like grid in the latitude-longitude direction, and then performs adaptive feature sampling on the cylinder map by deformable offset learning. Furthermore, to better integrate the features of the latitude-longitude plane and the spatial geometric features, we perform a multi-scale fusion of the extracted latitude-longitude features and spatial geometric features, and realize it through the aggregation of adjacent point features of different scales. In addition, a Sphere-wise Deformable point Convolution (SpDConv) operator is introduced to adaptively offset the sampling points in three-dimensional space by constructing a sphere grid structure, aiming at modeling the characteristics in the altitude direction. Experiments on existing popular benchmarks conclude that our TDConvs achieve the best segmentation performance, surpassing the existing state-of-the-art methods. The code is available on https://github.com/WingkeungM/TDConvs.
keywords:
[…]
1 Introduction
@@
In recent years, with the rapid advancement of deep learning technology and the explosive growth of remote sensing satellite data, the interpretation methods of remote sensing data have achieved unprecedented development. Especially the research on 3D remote sensing data has become increasingly popular. As the main sensor for collecting point cloud data in the field of remote sensing, airborne laser scanning (ALS) technology, has gradually entered the field of vision of researchers. Among them, ALS point cloud segmentation has received a lot of attention.
@@
Fig. 1.: ALS point clouds are acquired by the aircraft, therefore, these point clouds are strictly arranged according to the latitude and longitude direction and the altitude direction. (a) Latitude and Longitude Distribution. The instance objects in the large-scale point cloud are strictly arranged in fixed latitude and longitude positions. (b) Altitude Distribution. Different instance objects have different altitudes, which is one of the inherent characteristics of large-scale city-level point clouds.
@@
2 Related Work
2.1 Classical Machine Learning-Based Methods
@@
2.2 Point Cloud Segmentation Based on Deep Learning
The rapid development of deep learning has promoted the replacement of manual feature design methods by automatic point cloud feature extraction[51, 47, 39]. Among them, point-based methods are the most favored by researchers among many methods. The point-based methods[39, 40, 19, 60, 12, 56] refer to directly processing the point cloud without any data preprocessing operation, and then achieving the classification and segmentation tasks of the input point set by learning the deep semantic features of the point set. It mainly includes MLP-based methods[39, 40, 29, 52], graph-based methods[49, 43, 46], and point convolution-based methods[26, 24, 53, 47, 15]. MLP-based methods[39, 40, 29, 52] are the earlier deep convolutional neural network methods for processing point cloud data. PointNet[39] is the pioneering work of this type of method. It first uses multiple multi-layer perceptrons with shared weights to learn point-by-point high-dimensional features, and then designs a symmetric function based on the maximum pooling function to solve the disorder problem of point cloud. PointNet++[40] models local relationships through the three processes of sampling, grouping, and feature learning inside the point cloud. Point sets are the discrete structure, and there is a certain connection relationship between the points. This is a natural graph structure data form. Therefore, some researchers have introduced graph structure modeling into point cloud learning and achieved excellent results. Landrieu et al.[21] earlier proposed a point cloud convolutional neural network SPGraph based on graph convolution, which used graph convolution to learn high-level semantic features of the graph structure, and ultimately achieved point-by-point semantic segmentation. Although traditional convolution has been successful in feature extraction of 2D raster image data, it cannot be directly transferred to point cloud processing problems due to the unique properties of point cloud data. Therefore, Li et al.[26] tried to weight the features of the input point set and use a multi-layer perceptron to learn the X transformation to achieve point-by-point weighted summation in the neighborhood. In addition, Li et al.[24] proposed a geometric attention network composed of geometry-aware convolution, dense hierarchical architecture, and elevation attention module, and trained in an end-to-end manner to solve the three problems (geometric instances, extreme scale changes, and large differences in elevation) of remote sensing point clouds.
@@ 2.2 Point Cloud Segmentation Based on Deep Learning
Fig. 2.: Overall Architecture. Our TDConvs adopt the popular Encoder-Decoder structure as the basic component. Each feature extraction layer in the encoder part is the proposed Cylinder-wise Deformable point Convolution (CyDConv), and each layer in the decoder part is a feature forward propagation layer. The encoder and decoder are connected as feature aggregation layers with our proposed Sphere-wise Deformable point Convolution (SpDConv) at the highest resolution layer. The dimension of input point sets is $\rm N\times M$, where $\rm M$ is the dimension of point features. The dimension of output point sets is $\rm N\times C$, where $\rm C$ is the number of output categories.
2.3 Geospatial Modeling in Point Cloud Segmentation
@@
3 Proposed Method
3.1 Method Overview
@@
As shown in Fig. 2.2, the flowchart of our entire framework is given. Given the sampled point set, the encoder and decoder are executed sequentially. In addition, there is a skip-connection layer between each layer of the encoder and decoder for feature fusion. We embed the learning of longitude and latitude information into Cylinder-wise Deformable point Convolution (CyDConv), and the learning of altitude information into Sphere-wise Deformable point Convolution (SpDConv). CyDConv is performed in each feature extraction layer of the encoder, and SpDConv is performed in the highest resolution feature aggregation layer. The decoder is rooted in the feature propagation, which is proposed in PointNet++[40].
3.2 Cylinder-wise Deformable Point Convolution
@@
Cylindricization and Map Generation. Specifically, given the input point set $\mathbf{P}\in\mathbb{R}^{\rm N\times 3}$ and its feature $\mathbf{F}\in\mathbb{R}^{\rm N\times M}$, we first normalize the coordinates of the input point set $\mathbf{P}$, and then convert the point cloud into a cylinder structure at the top view, called cylindricization. Mathematically, we first construct the center points through the grid arrangement as shown in the bottom of Fig. 3.2, which matches each cylinder of the top view, as
@@
Fig. 3.: The architecture diagram of our proposed Cylinder-wise Deformable point Convolution (CyDConv). ‘LBR’ means the Linear-BN-ReLU operation. The parts from top to bottom in the figure represent Multi-Nearest Neighbor Feature Learning (MFL), Map Feature Sampling and Aggregation, Map Offset Learning, and Cylindricization and Map Generation processes, respectively.
Map Offset Learning. As shown in Fig. 3.2, to model the relationship within point sets in the latitude and longitude directions, we learn the offset through the point coordinates and point features, as
@@
Multi-Nearest Neighbor Feature Learning (MFL). From the perspective of the cylinder map, we have realized the modeling and learning of latitude and longitude features. For the modeling of geometric features inside point sets, we propose a Multi-nearest neighbor Feature Learning method called MFL. Specifically, we learn point cloud features by constructing neighboring points of different scales and then perform feature aggregation to realize geometric modeling of point cloud features of different scales, as shown in the top of Fig. 3.2.
Fig. 4.: The architecture diagram of the proposed Sphere-wise Deformable point Convolution (SpDConv). ‘T’ is a switch for whether to use SpDConv, which determines which connection path between the encoder and decoder uses SpDConv. Only the highest resolution connection path is selected to embed SpDConv in our network.
@@
3.3 Sphere-wise Deformable Point Convolution
@@
Spheroidization and Volume Generation. As shown in Fig. 2.2, our proposed SpDConv is embedded in the feature propagation layer between the encoder and the decoder. Since the highest-resolution point set has the most comprehensive altitude features, our SpDConv only selects the highest-resolution feature aggregation layer for embedding.
From Fig. 3.2, given the highest resolution point set $\mathbf{P}\in\mathbb{R}^{\rm N\times 3}$ and point feature $\mathbf{F}\in\mathbb{R}^{\rm N\times M}$. We first choose whether to close switch ‘T’. When switch ‘T’ is closed, SpDConv starts to perform operations. If the switch ‘T’ is closed, unlike CyDConv, we first build a three-dimensional grid structure. The size of the grid structure is $\rm H_{s}$,$\rm W_{s}$, and $\rm Z_{s}$, which is expressed by the formula
@@ 3.4 Loss Function
4 Experiments
4.1 Experimental Setting
4.1.1 Datasets
We selected the ISPRS Vaihingen 3D[7, 41, 37], LASDU[55], and DFC2019[22] datasets for experimental verification.
@@ 4.1.1 Datasets
DFC2019: The DFC2019 dataset[22] is a widely used ALS large-scale point cloud data benchmark in the remote sensing field. The data benchmark was collected from two areas in Jacksonville and Omaha, USA, covering an area of about 100 square kilometers. Each point in the dataset contains five main features, namely xyz coordinates, intensity, and return number. In addition, as a semantic segmentation dataset, its categories consist of about 200 million points in 5 categories.
4.1.2 Implementation Details
In the experiment, we adopted the PyTorch framework to complete our method. Specifically, for the data preprocessing process, we divide the point cloud of the datasets into patches according to a certain distance step (ISPRS Vaihingen 3D: $30m\times 30m$, LASDU: $50m\times 50m$, and DFC2019: $70m\times 70m$). For the training process, we randomly sample 4096 points from each patch as the input point sets of our TDConvs. In addition, the batch size and learning rate are set as 4 and 0.0002, respectively. We trained for 500 epochs on a single A40 GPU. What’s more, we adopt the Adam optimizer to minimize our target loss function. During testing, we take all points as input and evaluate the output results. The weights of our loss function are set to $\{1.0,2.0,2.0,2.0,2.0\}$ for the ISPRS Vaihingen 3D dataset, $\{1.0,5.0,5.0,5.0,5.0\}$ for LASDU dataset, and $\{1.0,5.0,5.0,5.0,5.0\}$ for DFC2019 dataset.
In the input process of the network, the ISPRS Vaihingen 3D, LASDU, and DFC2019 datasets all adopt coordinate information and feature information as input features of the point set. The feature information of the ISPRS Vaihingen 3D dataset contains normalized XYZ coordinates and intensity information. The feature information of the LASDU dataset contains normalized XYZ coordinates, RGB normalization information, and intensity information. The feature information of the DFC2019 dataset contains normalized XYZ coordinates and intensity information.
4.1.3 Evaluation Metrics
$\rm Precision=\frac{TP}{TP+FP}$ (20)
$\rm Recall=\frac{TP}{TP+FN}$ (21)
$\rm OA=\frac{TP+TN}{TP+TN+FP+FN}$ (22)
[…]
What’s more, mF1 are the averages of F1 for all categories.
4.2 Performance Analysis
4.2.1 Quantitative Analysis
We present the segmentation performance of our TDConvs on three datasets, ISPRS Vaihingen 3D, LASDU, and DFC2019, respectively.
Table 1: Performance of our TDConvs and other methods on the ISPRS Vaihingen 3D dataset. ‘vision’ means vision-based methods. ‘Remote Sensing’ represents remote sensing-based methods. The first 9 columns of the table represent the F1 score of each category. OA: overall accuracy. mF1: mean F1 score. Bold indicates the first performance, and underline indicates the second performance.
@@ 4.2.1 Quantitative Analysis
Table 2: Performance comparison of our TDConvs with other methods on the LASDU dataset. Bold indicates the first performance and underline indicates the second performance.
@@
ISPRS Vaihingen 3D: As shown in Table 1, we present the performance comparison of our TDConvs with other point-based methods on the ISPRS Vaihingen 3D dataset. In general, our TDConvs achieve 84.5% on OA and 73.4% on mF1, respectively, surpassing all other remote sensing-based and vision-based methods, which strongly proves the superiority of our TDConvs. Specifically, our TDConvs outperforms the popular RFFS-Net[32] by 2.4% on OA and 1.8% on mF1. OA and mF1 evaluate the overall accuracy and category accuracy of the model, respectively. Obviously, our TDConvs achieves the best performance for both OA and mF1, and significantly outperforms the SOTA method. For specific categories, our TDConvs achieves balanced performance on each category with small variance in performance, and there are no categories with extremely low performance. This is because ALS point clouds are strictly arranged according to latitude, longitude, and altitude. Our TDConvs can achieve balanced segmentation performance. In particular, for categories of car, fence, and facade, our TDConvs achieves the best or second segmentation performance, which strongly demonstrates the superiority of our approach.
Table 3: Performance comparison of our TDConvs with other methods on the DFC2019 dataset. Bold indicates the first performance and underline indicates the second performance.
Method Ground Trees Buildings Water Bridge OA mF1
PointNet++ [40] 98.30 95.80 79.70 4.40 7.30 92.70 57.10
PointSIFT [17] 98.60 97.00 85.50 46.40 60.40 94.00 77.60
PointCNN [26] 98.70 97.20 84.90 44.10 65.30 93.80 78.00
KPConv [47] 98.40 94.20 87.40 43.00 77.50 94.50 80.10
DGCNN [51] 97.88 93.22 90.37 88.23 54.39 95.08 84.82
D-FCN [52] 99.10 98.10 89.90 45.00 73.00 95.60 81.00
DANCE-NET [25] 99.10 93.90 87.00 58.30 83.90 96.80 84.40
PointConv [53] 97.33 95.82 93.63 74.50 69.24 95.32 86.10
RFFS-Net [32] 96.61 96.10 88.69 77.84 80.97 94.31 88.04
LGGAA [18] 98.90 96.10 90.20 41.60 83.70 94.80 81.40
VD-LAB [23] 98.04 95.53 92.07 85.51 71.45 95.93 88.52
IPCONV [62] 98.80 97.50 92.90 92.10 58.20 97.10 87.90
LGENet [27] 99.30 98.30 92.80 47.40 79.10 98.40 83.40
DA-Net [61] 99.30 97.60 92.70 41.60 85.10 98.30 83.30
RRDAN [59] 99.10 98.10 95.80 62.80 82.30 98.10 87.60
TDConvs (ours) 99.17 97.64 96.84 94.41 86.80 98.24 94.97
Fig. 5.: Segmentation results of the proposed TDConvs on the ISPRS Vaihingen 3D test set. The results in the box are local magnifications of the large area.
@@
LASDU: The comparison on the LASDU dataset of our TDConvs and other methods is given in Table 2. From the table, it can be concluded that our proposed TDConvs achieves the best segmentation performance, reaching an mF1 score of 77.85%. In addition, on the OA indicator, our method is comparable to the optimal MCFN, but on the mF1 indicator, the TDConvs we proposed is 0.85% higher than MCFN[58], which is a big improvement in ALS point cloud segmentation. Compared with the popular RFFS-Net[32], our proposed TDConvs is 0.49% and 0.16% higher in OA and mF1 indicators respectively, which also reflects the superiority of the method in this paper. Compared with OA, mF1 can better express the performance of point cloud segmentation because it takes into account the performance of each category, which is particularly obvious in datasets with imbalanced categories. Therefore, in the field of segmentation, we focus more on the performance of mF1, and on the LASDU dataset, our TDConvs achieve the best mF1 performance, which strongly proves the superior performance of our TDConvs in the field of point cloud segmentation. As shown in Table 2, in specific categories, although our method does not achieve the best segmentation performance, the performance of each category is balanced. This has obvious advantages over IPCONV[62], which has large segmentation differences in categories. This is why our method can achieve the best mean F1 performance.
DFC2019: As shown in Table 3, we give the performance comparisons of existing methods and our proposed TDConvs on the DFC2019 dataset. Obviously, our TDCOnvs achieves the best mF1 performance of 94.97%, which is a high performance in the field of large-scale point cloud semantic segmentation. In addition, our method is 6.45% higher than the existing SOTA method VD-LAB[23], which is a huge improvement. Through careful comparison, our TDConvs achieves the best performance in the Buildings, Water, and Bridge categories, the second best performance in the Ground category, and comparable performance to SOTA methods in the Trees category. This is because the TDConvs we proposed can explicitly model the latitude, longitude, and altitude information, which plays an important role in ensuring the balance of performance in each category. This is also the reason why our method can achieve superior performance in each category.
4.2.2 Qualitative Analysis